# Split-and-Denoise: Protect Large Language Model Inference with Local Differential Privacy

### v3

@@ 2 Prior Works
With the advent of LLMs, privacy leakage has emerged as a crucial concern. Existing literature predominantly focuses on privacy protection throughout the entire training process, encompassing pre-training (Hoory et al., 2021), fine-tuning (Huang et al., 2020; Kerrigan et al., 2020; Yu et al., 2021; Lukas et al., 2023; Shen et al., 2023), and prompt-tuning phases (Duan et al., 2023; Li et al., 2023). Yet, there is a notable dearth of research that addresses local privacy during the inference phase with a fully frozen LLM. This scenario, which prohibits alterations to the model’s structure and parameters, is particularly complex. Nonetheless, it holds significance in black-box API access contexts, especially for proprietary models like GPT-4. An intuitive approach involves anonymizing sensitive terms prior to LLM input and subsequently restoring them post-output (Kan et al., 2023; Chen et al., 2023). However, this method, while effective for obfuscating specific entities, falls short in concealing other linguistic elements, including verbs and non-named entities. Such a limitation compromises full privacy and is unsuitable for tasks necessitating exact semantic interpretation of the altered entities, such as knowledge retrieval and text continuation (Chen et al., 2023). An alternative strategy might entail privatizing the input at token representations or intermediate layer levels. Qu et al. (2021b) investigates the utility and privacy tradeoff for privacy-preserving finetuing, involving text-to-text privatization  (Feyisetan et al., 2019; Qu et al., 2021a) and token embedding privatizations, while the two techniques could be adapted to private LLM inference. Privacy-Preserving Prompt Tuning (RAPT) (Li et al., 2023) employs text-text privatization to conduct prompt tuning and inference with local differential privacy. The authors propose a reconstruction head during prompt tuning to enhance the utility. Another direction employs homomorphic encyption (HE) to conduct private transformer inference such as Privacy-Computing Friendly Transformers (PCFT) and The-x (Liu & Liu, 2023; Chen et al., 2022), but the significant overhead renders it impractical for implementation in LLM.
@@
Split learning is a privacy-preserving approach in distributed learning, where each client trains a segment of a deep network up to a designated ”cut layer.” The outputs at this layer are then forwarded to the server side, which completes the training without accessing the client’s raw data. This approach facilitates forward and backward propagation without sharing raw data, ensuring the client-side local privacy  (Gupta & Raskar, 2018; Vepakomma et al., 2018). Vepakomma et al shows that split learning surpasses federated learning and large batch synchronous SGD in achieving superior accuracy with significantly reduced client-side computational demands (Gupta & Raskar, 2018). Singh et al further validate its efficacy across broader experimental contexts, demonstrating that an increase in the number of clients or model dimensions gives split learning an edge over federated learning (Singh et al., 2019). The advantage in its computational efficiency renders it suitable for LLM local privacy setting, where the client side executes minimal computational tasks, such as noising and denoising operations at specific segmented layers, to ensure privacy at reduced computational expenses. Meanwhile, the server handles the bulk of the model’s layers. Our research serves as an initial endeavor to integrate split learning with LLM privacy concerns.
@@ 5 Discussion and Future Work
Reduce user computation cost: local denoising constitutes a major component of user’s computation overhead. We observe that the size of denoise model, and thus the user computation cost, scale with the underlying LLM. For those users with limited computation resource, it’s crucial to design a lightweight denoise mechanism with minimal computation cost.
Sequence-to-sequence (S2S) inference: it’s of great interest to extend our EaaS framework to S2S inference model. One important obstacle of the generalization is the noise amplification issue with S2S model. In particular, S2S relies on the auto-regressive mechanism, where the prediction of previous token is taken as an input to the next token. Therefore, the error from the previous prediction would exaggerate the deviation of the following tokens. A universal denoise model might be insufficient to correct the errors in the generated sequence.
@@ A.5.3 Training of Denoise Model
GPT2 GPT2 Small 5.8 hours on single A6000 ã week using several NVIDIA V100 GPUs
@@ A.12 Ablation Studies
Impact of server-side denoise model: to evaluate the impact of noise level awareness on denoising performance, we deploy the denoise model on server side using only privatized token representations as input. We train a transformer-based denoise model that outputs the denoised token representations with the objective to minimize MSE. The denoise model adopts the same hyperparameters specified in Appendix A.5. Figure 7 demonstrates that SnD significantly outperforms the server-side denoise method by over 10% in almost all cases. It’s also important to note that most AUCs of server-side denoise model fall below 0.55, suggesting the server’s incapacity to deduce private information about the user.

### v4

@@ 2 Prior Works
With the advent of LLMs, privacy leakage has emerged as a crucial concern. Existing literature predominantly focuses on privacy protection throughout the entire training process, encompassing pre-training (Hoory et al., 2021; Ding et al., 2024), fine-tuning (Huang et al., 2020; Kerrigan et al., 2020; Yu et al., 2021; Lukas et al., 2023; Shen et al., 2023; Ye et al., 2024), and prompt-tuning phases (Duan et al., 2023; Li et al., 2023). Yet, there is a notable dearth of research that addresses local privacy during the inference phase with a fully frozen LLM. This scenario, which prohibits alterations to the model’s structure and parameters, is particularly complex. Nonetheless, it holds significance in black-box API access contexts, especially for proprietary models like GPT-4. An intuitive approach involves anonymizing sensitive terms prior to LLM input and subsequently restoring them post-output (Kan et al., 2023; Chen et al., 2023). However, this method, while effective for obfuscating specific entities, falls short in concealing other linguistic elements, including verbs and non-named entities. Such a limitation compromises full privacy and is unsuitable for tasks necessitating exact semantic interpretation of the altered entities, such as knowledge retrieval and text continuation (Chen et al., 2023). An alternative strategy might entail privatizing the input at token representations or intermediate layer levels. Qu et al. (2021b) investigates the utility and privacy tradeoff for privacy-preserving finetuing, involving text-to-text privatization  (Feyisetan et al., 2019; Qu et al., 2021a) and token embedding privatizations, while the two techniques could be adapted to private LLM inference. Privacy-Preserving Prompt Tuning (RAPT) (Li et al., 2023) employs text-text privatization to conduct prompt tuning and inference with local differential privacy. The authors propose a reconstruction head during prompt tuning to enhance the utility. Another direction employs homomorphic encyption (HE) to conduct private transformer inference such as Privacy-Computing Friendly Transformers (PCFT) and The-x (Liu & Liu, 2023; Chen et al., 2022), but the significant overhead renders it impractical for implementation in LLM.
@@
Split learning is a privacy-preserving approach in distributed learning, where each client trains a segment of a deep network up to a designated ”cut layer.” The outputs at this layer are then forwarded to the server side, which completes the training without accessing the client’s raw data. This approach facilitates forward and backward propagation without sharing raw data, ensuring the client-side local privacy  (Gupta & Raskar, 2018; Vepakomma et al., 2018). Vepakomma et al. shows that split learning surpasses federated learning and large batch synchronous SGD in achieving superior accuracy with significantly reduced client-side computational demands (Gupta & Raskar, 2018). Singh et al. further validate its efficacy across broader experimental contexts, demonstrating that an increase in the number of clients or model dimensions gives split learning an edge over federated learning (Singh et al., 2019). Ding et al. applies the idea of split learning to the computer vision task (Ding et al., 2022). The advantage in its computational efficiency renders it suitable for LLM local privacy setting, where the client side executes minimal computational tasks, such as noising and denoising operations at specific segmented layers, to ensure privacy at reduced computational expenses. Meanwhile, the server handles the bulk of the model’s layers. Our research serves as an initial endeavor to integrate split learning with LLM privacy concerns.
@@ 5 Discussion and Future Work
Reduce user computation cost: local denoising constitutes a major component of user’s computation overhead. We observe that the size of denoise model, and thus the user computation cost, scale with the underlying LLM. For those users with limited computation resource, it is crucial to design a lightweight denoise mechanism with minimal computation cost.
Sequence-to-sequence (S2S) inference: it is of great interest to extend our EaaS framework to S2S inference model. One important obstacle of the generalization is the noise amplification issue with S2S model. In particular, S2S relies on the auto-regressive mechanism, where the prediction of previous token is taken as an input to the next token. Therefore, the error from the previous prediction would exaggerate the deviation of the following tokens. A universal denoise model might be insufficient to correct the errors in the generated sequence.
@@ A.5.3 Training of Denoise Model
GPT2 GPT2 Small 5.8 hours on single A6000 A week using several NVIDIA V100 GPUs
@@ A.12 Ablation Studies
Impact of server-side denoise model: to evaluate the impact of noise level awareness on denoising performance, we deploy the denoise model on server side using only privatized token representations as input. We train a transformer-based denoise model that outputs the denoised token representations with the objective to minimize MSE. The denoise model adopts the same hyperparameters specified in Appendix A.5. Figure 7 demonstrates that SnD significantly outperforms the server-side denoise method by over 10% in almost all cases. It is also important to note that most AUCs of server-side denoise model fall below 0.55, suggesting the server’s incapacity to deduce private information about the user.